OpenStudy (anonymous):

Just want to make sure I can do this Got the integral of e^x * e^x dx. I make u=e^x, and du has to be e^x as well. So I just use one for each, and end up with integral of u du.

OpenStudy (astrophysics):

\[\int\limits e^x e^x dx\]

OpenStudy (astrophysics):

?

OpenStudy (astrophysics):

\[\int\limits\limits e^x e^x dx = \int\limits e^{2x} dx\]

OpenStudy (astrophysics):

u = 2x yadi yadi yada

Parth (parthkohli):

or sub \(u = e^x\)

Parth (parthkohli):

oh @DurableToaster has written that already oops

OpenStudy (astrophysics):

yeah

Parth (parthkohli):

why not?

OpenStudy (astrophysics):

i dont talk to kids with cp

Parth (parthkohli):

hey I don't have CP just chicken pox

Parth (parthkohli):

but yes, not talking to people with CP is the way to go

OpenStudy (astrophysics):

thanks

OpenStudy (anonymous):

That's strange. Couldn't I just take the whole e^x as u? Cause d/dx e^x is e^x, u=du, so I could substitute the e^xs for different things, getting e^x^2 * 1/2.... The Same as you got. Sweet!

OpenStudy (astrophysics):

Yes you can